OpenStudy (arbershabani97):

Help!! 2^x=.4 how to solve without a calc

OpenStudy (anonymous):

well you square both sides which cancels out the x and you get X=2

OpenStudy (arbershabani97):

ooops very sorry i wrote it wrong 2^x=.4

OpenStudy (anonymous):

oh haha

OpenStudy (anonymous):

okay! gimme a sec.

OpenStudy (arbershabani97):

okay :p

OpenStudy (anonymous):

since you cant make the base = you have to use log

OpenStudy (arbershabani97):

but in class i learned only how to solve without calculator, and i will learn it in the next sections, may I know how to do that thanks

OpenStudy (anonymous):

take this problem \[7^x=12\] \[\ln 7^x=\ln 12\] \[x \ln 7= \ln 12\] \[x=\frac{ \ln 12 }{ \ln 7 }\] \[\approx 1.277\] this is straight out of my algebra book.

OpenStudy (arbershabani97):

thank you very much